### v1

@@ 5.4.2 Preprocessing
  • Text Embedding Extraction: We used the language foundational model Llama 2 with its 7 billion parameter configuration to convert the text in the clinical notes into vector embeddings, enabling a nuanced understanding of clinical notes in a lower dimensional and computationally efficient representation.
@@ 6 Discussion
The framework proposed in this paper presents a novel approach for multimodal data fusion centered on the use of embeddings, foundational models, and data mining techniques. It addresses an efficient process for extracting knowledge from diverse data modalities while considering resource constraints. In this section, we discuss the advantages and implications of our approach, contrast it with existing models, and discuss potential limitations.
@@ 6.1 Advantages of the Proposed Framework
  • Dimensionality Reduction with Embeddings: Using foundational models and embeddings introduces an efficient solution for high-dimensional modalities like text and images. This method reduces computational costs for model training and data storage, making it a practical choice for resource-constrained environments.
@@
  • Bias mitigation: Our framework incorporates a dedicated level (Level 7 - Bias Assessment) to systematically address and mitigate bias throughout the entire DF-DM model. By recognizing that bias can manifest at different stages, we provide some strategies for each aspect. This proactive approach involves techniques such as diverse data collection, bias detection, preprocessing, analysis of foundational models, causal inference, bias awareness training, and audits or peer review.
@@ 6.2 Comparison with Existing Models
While traditional data fusion models, such as the DFGI model, offer a structured approach, they often need more flexibility and efficiency to handle diverse data modalities, particularly in resource-limited settings. Our proposed framework bridges these gaps by drawing inspiration from the well-established DFGI model and integrating the flexibility and modern data understanding elements of the CRISP-DM model and the most up-to-date DL techniques. The introduction of foundational models and embeddings further enhances the efficiency and effectiveness of the proposed model.
@@ 6.4.2 Use case 2: Domestic Violence Prediction Using Open Data
Notably, the use of self-supervised learning in domain-specific scenarios, such as our Resnet 50 V2 VAE model, outperformed foundational computer vision models like Dino V2. This highlights the importance of domain specific knowledge for data analysis and model training.
@@ 6.4.3 Use case 3: MIMIC CXR Clinical notes and x-ray images
The MIMIC-CXR dataset presented an evaluation of the DF-DM model’s efficacy across both clinical and fairness tasks, specifically disease diagnosis and sex prediction. This dual-focused analysis underscores the versatility of our model, not only in tackling complex medical diagnostic challenges but also in addressing critical issues of bias within AI-driven healthcare solutions. The implementation of our model on this extensive dataset, encompassing a broad spectrum of chest radiographs paired with clinical notes, highlights the model’s capability to process and derive insights from multimodal data effectively. This ability is paramount in the context of healthcare, where the integration of various data types can significantly enhance diagnostic accuracy and patient outcomes. By achieving superior, performance metrics against state-of-the-art models in these tasks, our model demonstrates its potential to serve as a valuable tool in clinical settings. The inclusion of sex prediction as a fairness task further enriches our analysis, providing a platform to detect and mitigate potential biases inherent in machine learning models. Such considerations are crucial in ensuring that AI-driven healthcare solutions promote equity and do not perpetuate existing disparities. The successful application of the DF-DM model in this use case not only validates its conceptual framework but also accentuates its practical utility in real-world scenarios. By navigating the challenges presented by the high-dimensional and heterogeneous nature of the MIMIC-CXR dataset, our model underscores the importance of leveraging embeddings and foundational models to streamline data processing and enhance computational efficiency.
@@ 6.5 Limitations and Future Directions
It is essential to acknowledge that our framework, while addressing many challenges in multimodal data fusion, has limitations. The effectiveness of our approach depends on the availability of foundational models, which may not always be accessible in specific domains. Additionally, the human-in-the-loop approach introduced in Level 5 may require additional resources and time. As a future direction, research can focus on optimizing the framework for scenarios where foundational models are not readily available, exploring the development of more efficient techniques for dimensionality reduction, and further refining the interaction between human experts and the model to minimize resource requirements.
@@ 7 Conclusion
In this paper, we introduced the Data Fusion for Data Mining (DF-DM) model, a groundbreaking process model that not only leverages the power of embeddings and foundational models for efficient data fusion but also introduces a novel mutual information based dense fusion model. By optimizing for mutual information, our model ensures that the fused data representation captures the most relevant shared information between modalities while ensuring modality-specific expressiveness, significantly improving the cohesion and efficacy of the fusion process. Our approach enhances efficiency, flexibility, and data understanding while considering resource constraints. By bridging the gaps in existing models, we provide a pathway for improved data fusion in a wide range of applications. We also provided a validation of the model showcasing three healthcare scenarios proving the effectiveness of the approach in healthcare settings. The proposed DF-DM and mutual embedding information alignment model offer a foundational approach to multimodal data fusion, addressing the challenges of efficiency, flexibility, and data understanding. However, it is essential to acknowledge and actively mitigate biases that may arise in the complex process of integrating diverse data modalities. Including Level 7 - Bias Assessment and the outlined strategies reflect our commitment to promoting fairness, equity, and reliability in the fusion of multimodal data. As the volume and complexity of multimodal data continue to grow, this framework offers a promising solution that combines the best of traditional data fusion models with the most used techniques in data mining and deep learning with embeddings and foundational models. This research provides a way for more efficient and resource-aware data fusion processes, contributing to advancements in healthcare, environmental sciences, and beyond.

### v2

@@ 5.4.2 Preprocessing
  • Text Embedding Extraction: We used the language foundation model Llama 2 with its 7 billion parameter configuration to convert the text in the clinical notes into vector embeddings, enabling a nuanced understanding of clinical notes in a lower dimensional and computationally efficient representation.
@@ 6 Discussion
The framework proposed in this paper presents a novel approach for multimodal data fusion centered on the use of embeddings, foundation models, and data mining techniques. It addresses an efficient process for extracting knowledge from diverse data modalities while considering resource constraints. In this section, we discuss the advantages and implications of our approach, contrast it with existing models, and discuss potential limitations.
@@ 6.1 Advantages of the Proposed Framework
  • Dimensionality Reduction with Embeddings: Using foundation models and embeddings introduces an efficient solution for high-dimensional modalities like text and images. This method reduces computational costs for model training and data storage, making it a practical choice for resource-constrained environments.
@@
  • Bias mitigation: Our framework incorporates a dedicated level (Level 7 - Bias Assessment) to systematically address and mitigate bias throughout the entire DF-DM model. By recognizing that bias can manifest at different stages, we provide some strategies for each aspect. This proactive approach involves techniques such as diverse data collection, bias detection, preprocessing, analysis of foundation models, causal inference, bias awareness training, and audits or peer review.
@@ 6.2 Comparison with Existing Models
While traditional data fusion models, such as the DFGI model, offer a structured approach, they often need more flexibility and efficiency to handle diverse data modalities, particularly in resource-limited settings. Our proposed framework bridges these gaps by drawing inspiration from the well-established DFGI model and integrating the flexibility and modern data understanding elements of the CRISP-DM model and the most up-to-date DL techniques. The introduction of foundation models and embeddings further enhances the efficiency and effectiveness of the proposed model.
@@ 6.4.2 Use case 2: Domestic Violence Prediction Using Open Data
Notably, the use of self-supervised learning in domain-specific scenarios, such as our Resnet 50 V2 VAE model, outperformed foundation computer vision models like Dino V2. This highlights the importance of domain specific knowledge for data analysis and model training.
@@ 6.4.3 Use case 3: MIMIC CXR Clinical notes and x-ray images
The MIMIC-CXR dataset presented an evaluation of the DF-DM model’s efficacy across both clinical and fairness tasks, specifically disease diagnosis and sex prediction. This dual-focused analysis underscores the versatility of our model, not only in tackling complex medical diagnostic challenges but also in addressing critical issues of bias within AI-driven healthcare solutions. The implementation of our model on this extensive dataset, encompassing a broad spectrum of chest radiographs paired with clinical notes, highlights the model’s capability to process and derive insights from multimodal data effectively. This ability is paramount in the context of healthcare, where the integration of various data types can significantly enhance diagnostic accuracy and patient outcomes. By achieving superior, performance metrics against state-of-the-art models in these tasks, our model demonstrates its potential to serve as a valuable tool in clinical settings. The inclusion of sex prediction as a fairness task further enriches our analysis, providing a platform to detect and mitigate potential biases inherent in machine learning models. Such considerations are crucial in ensuring that AI-driven healthcare solutions promote equity and do not perpetuate existing disparities. The successful application of the DF-DM model in this use case not only validates its conceptual framework but also accentuates its practical utility in real-world scenarios. By navigating the challenges presented by the high-dimensional and heterogeneous nature of the MIMIC-CXR dataset, our model underscores the importance of leveraging embeddings and foundation models to streamline data processing and enhance computational efficiency.
@@ 6.5 Limitations and Future Directions
It is essential to acknowledge that our framework, while addressing many challenges in multimodal data fusion, has limitations. The effectiveness of our approach depends on the availability of foundation models, which may not always be accessible in specific domains. Additionally, the human-in-the-loop approach introduced in Level 5 may require additional resources and time. As a future direction, research can focus on optimizing the framework for scenarios where foundation models are not readily available, exploring the development of more efficient techniques for dimensionality reduction, and further refining the interaction between human experts and the model to minimize resource requirements.
@@ 7 Conclusion
In this paper, we introduced the Data Fusion for Data Mining (DF-DM) model, a groundbreaking process model that not only leverages the power of embeddings and foundation models for efficient data fusion but also introduces a novel mutual information based dense fusion model. By optimizing for mutual information, our model ensures that the fused data representation captures the most relevant shared information between modalities while ensuring modality-specific expressiveness, significantly improving the cohesion and efficacy of the fusion process. Our approach enhances efficiency, flexibility, and data understanding while considering resource constraints. By bridging the gaps in existing models, we provide a pathway for improved data fusion in a wide range of applications. We also provided a validation of the model showcasing three healthcare scenarios proving the effectiveness of the approach in healthcare settings. The proposed DF-DM and mutual embedding information alignment model offer a foundational approach to multimodal data fusion, addressing the challenges of efficiency, flexibility, and data understanding. However, it is essential to acknowledge and actively mitigate biases that may arise in the complex process of integrating diverse data modalities. Including Level 7 - Bias Assessment and the outlined strategies reflect our commitment to promoting fairness, equity, and reliability in the fusion of multimodal data. As the volume and complexity of multimodal data continue to grow, this framework offers a promising solution that combines the best of traditional data fusion models with the most used techniques in data mining and deep learning with embeddings and foundation models. This research provides a way for more efficient and resource-aware data fusion processes, contributing to advancements in healthcare, environmental sciences, and beyond.